OpenStudy (anonymous):

The vectors a-5b and a-b are perpendicular. If a and b are unit vectors, then determine a (dot) b.

OpenStudy (anonymous):

You have vectors \[v_1=\vec{a}-5\vec{b}\] and \[v_2=\vec{a}-\vec{b}\] they are perpendicular, so their dot product is...?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

but the answer is 1

OpenStudy (anonymous):

v1.v2=0 \[a ^{2}=1,b ^{2}=1,find a.b\]

OpenStudy (anonymous):

@surjithayer I don't get it...sorry

OpenStudy (anonymous):

@amistre64 please help

OpenStudy (anonymous):

Does this question have an answer? If you take the dot product of the two vectors which should be perpendicular, you have an expression for the ratio of the length of the unit vectors a and b. They are not 1 however...am I missing something?

OpenStudy (amistre64):

since they are being very general; lets go ahead and define some parts that would be useful to us: a = (1,0), and let b = (x,y) a-5b u = (1-5x,0-5y) v = ( 1-x , 0-y) then we can establish the dot product as zero

OpenStudy (anonymous):

@TimSmit the answer is 1, but I don't understand how they get that

OpenStudy (amistre64):

the dot product should give us the equation of a circle ...

OpenStudy (amistre64):

or an ellipse, maybe a hyperbola ..... something conic fer sure :) u = (1-5x,0-5y) v = ( 1-x , 0-y) --------------- (5x^2-6x+1) + (5y^2-6y) = 0 5(x^2-6x/5)+1 + 5(y^2 - 6y/5) = 0 (x^2-6x/5) +1/5 + (y^2 - 6y/5) = 0 (x^2-6x/5+36/100) - 36/100 +1/5 + (y^2 - 6y/5 + 36/100) -36/100 = 0 (x-6/10)^2 + (y-6/10)^2 + (20 - 72)/100 = 0 (x-6/10)^2 + (y-6/10)^2 = 27/50 yay!! its a circle

OpenStudy (amistre64):

where was i going with this ?.....

OpenStudy (amistre64):

oh right .... sqrt(x^2+y^2) = 1; therefore say: y = sqrt(1-x^2) and solve for x

OpenStudy (amistre64):

(1,0) dot (x,y) = x

OpenStudy (anonymous):

I can't seem to solve it :) the dot product of the two perpendicular vectors is: \[||a||^2+5||b||^2=0\] If a and b are unit vectors then their lengths are 1, so 6=0?

OpenStudy (anonymous):

Is there any simpler explanation?

OpenStudy (amistre64):

probably, but i tend to have to go off of what i can piece together from years gone by :)

OpenStudy (anonymous):

Could you check your vectors eLg? are they really a-5b and a-b?

OpenStudy (anonymous):

yes they are right

OpenStudy (amistre64):

\[(x_a-5x_b,y_a-5y_b)\cdot (x_a-x_b,y_a-y_b)=0\] i do get a solution close to 1, approx 0.99986...

OpenStudy (anonymous):

Ah I was thinking too perpendicular for unit vectors while taking the dot product.

OpenStudy (amistre64):

since the question is general, i dont see a reason why we couldnt define a or b as 1,0; lets use b instead and see how that goes \[(x_a-5,y_a)\cdot (x_a-1,y_a)=0\] \[(x^2-6x+5)+(y^2)=0\] \[(x^2-6x+9)+(y^2)=4\] since sqrt(x^2+y^2)=1, let y = sqrt(1-x^2) \[(x^2-6x+9)+(1-x^2)=4\] \[-6x+9+1=4\] \[-6x=-6\] \[x=1\]

OpenStudy (amistre64):

therefore: (x,y) dot (1,0) = x = 1

OpenStudy (amistre64):

i most likely did some bad mathing the first go around :)

OpenStudy (anonymous):

oh ok,

OpenStudy (anonymous):

Working out your dot product amistre: \[(x_a-5x_b,y_a-5y_b)\cdot(x_a-x_b,y_a-y_b)=0\] \[(x_a-5x_b)*(x_a-x_b)+(y_a-5y_b)*(y_a-y_b)=0\] \[x_a^2-6x_ax_b+5x_b^2+y_a^2-6y_ay_b+5y_b^2=0\] As a and b are unit vectors: \[\sqrt{x_a^2+y_a^2}=1=x_a^2+y_a^2\] Then \[-6x_ax_b-6y_ay_b+6=0\] Now, choosing for a=(1,0) gives: \[-6x_b+6=0\] \[x_b=1\] As b is a unit vector, b=(1,0) So the dot product of a and b is 1 as a=b. Correct?

OpenStudy (amistre64):

id have to take a few aspirins to be sure :)

OpenStudy (anonymous):

because they are perpendicular so dot product is 0 \[\left( vectora-vectorb \right).\left( vectora-vector5b \right)=0\] \[or\left( vectora \right)^{2}-5\left( vectora \right).\left( vector5b \right)-\left( vectorb \right).\left( vectora \right)+5\left( vectorb \right)^{2}= 0\] \[1-6 \left( vectora \right).\left( vectorb \right)+5\left( 1 \right)= 0 \] \[6\left( vectora \right).\left( vectorb \right) =6\] \[( vector a).\left( vectorb \right) =\frac{ 6 }{6 }=1 \]

OpenStudy (amistre64):

\[(x^2_a-6x_ax_b+5x^2_b)+(y^2_a-6y_ay_b+5y^2_b)=0\] \[y_a=\sqrt{1-x_a^2~}~:~y_b=\sqrt{1-x_b^2~}\] \[x^2_a-6x_ax_b+5x^2_b+(\sqrt{1-x_a^2~})^2-6\sqrt{(1-x_a^2)(1-x_b^2)}+5(\sqrt{1-x_b^2~})^2=0\] \[x^2_a-6x_ax_b+5x^2_b+1-x_a^2-6\sqrt{(1-x_a^2)(1-x_b^2)}+5-5x_b^2=0\] \[-6x_ax_b-6\sqrt{(1-x_a^2)(1-x_b^2)}+6=0\] \[1=x_ax_b+\sqrt{(x_ax_b)^2-x_a^2-x^2_b+1}\] and as long as i dont mess it up along the way :)

OpenStudy (amistre64):

pfft, i think i should have kept the ys in there

OpenStudy (phi):

surji showed how do it easily \[ (a-5b)\cdot ( a-b) =0 \\ a^2 -5(a\cdot b) - a\cdot b +5 b^2 =0\] a, and b are unit length. simplify and combine like terms \[ 1 -6(a\cdot b) + 5 =0 \\ a\cdot b= 1\]

OpenStudy (amistre64):

\[-6x_ax_b-6\sqrt{(1-x_a^2)(1-x_b^2)}+6=0\] \[-6x_ax_b-6\sqrt{(1-x_a^2)}\sqrt{(1-x_b^2)}+6=0\] \[x_ax_b+\sqrt{(1-x_a^2)}\sqrt{(1-x_b^2)}=1\] \[x_ax_b+y_ay_b=1\] and therefore the the defintion of a.b

OpenStudy (anonymous):

correction i typed 5 extra in the third line.